Known Issues

Known Issues

Title Category Resolution Description Posted Updated
Rolling reboot of Oakley and Ruby clusters, starting from 8:30AM October 9, 2017 Batch, login, Oakley, Ruby Resolved

Updates on 1:00PM October 16, 2017: 

The rolling reboots of Oakley and Ruby are completed. 

... Read more
1 week 6 days ago 2 days 16 hours ago
Rolling reboot of Owens cluster, starting from 9AM September 11, 2017 Batch, Owens Resolved

Updates on 12:20PM September 25, 2017: 

The rolling reboot of Owens is completed. 

... Read more
1 month 1 week ago 3 weeks 2 days ago
Some issues remain after downtime Login Problems, Operations, Outage Resolved

15 July 2016, 5:00PM update: some additional issues we are facing

  • We are experiencing periodic hangs of the GPFS client file system software used with the new storage environment. We... Read more
1 year 3 months ago 4 weeks 1 day ago
OnDemand has NOT been working with external providers since 08/22 OnDemand Resolved

Updates on 9:40AM August 23, 2017: this issue has been resolved. 

>>>

Issue:

User can NOT login to OnDemand (ondemand.osc.edu)... Read more

1 month 3 weeks ago 1 month 3 weeks ago
PBS commands on Owens are not working Batch, Owens Resolved

Update posted on July 12, 2017 at 1:50PM:

We have fixed the problem with the batch management system on Owens and queues on Owens have been opened again for jobs.

... Read more

3 months 1 week ago 3 months 1 week ago
Rolling reboot of Owens cluster, starting from 9AM June 28, 2017 Owens Resolved

Update posted on July 7, 2017 at 2:00PM:

Rolling reboot of login and compute nodes of Owens cluster is completed. 

... Read more
3 months 3 weeks ago 3 months 1 week ago
Systemic Problem on Cluster Computing service Operations Resolved

4:20PM 6/23/2017 Update: All HPC systems are back in production. This outage may cause failures of users' jobs. We'll update the community as more is known. 

... Read more
4 months 1 day ago 3 months 3 weeks ago
my.osc.edu is NOT available Account Management Resolved

my.osc.edu has not been fully restored after yesterday's downtime. You can change your password, but you will not be able to use the new password on my.osc.edu. The updated password will work to... Read more

4 months 3 weeks ago 4 months 3 weeks ago
"pbsdcp" is not working on Oakley Oakley Resolved

12:35PM 5/24/2017 Update: pbsdcp   has been fixed on Oakley.

pbsdcp   is not working on Oakley and returns a missing library error as below:... Read more

4 months 3 weeks ago 4 months 3 weeks ago
Issue with GPFS on Owens since April 14, 2017 Batch, filesystem, Owens Resolved

3:10PM 4/18/2017 Update: Rolling reboots on Owens have started to address this GPFS issue. 

We have had issues with GPFS mounts on Owens Cluster since Friday afternoon,... Read more

6 months 4 days ago 5 months 2 weeks ago

Pages